Two events A and B are such that, they are independent and P(A)=x and P(B)=x+0.2 and P(A n B )=0.15. Find;


i) the value of x.

ii) P(A U B) and P( Ac / Bc ), where Ac, Bc are the complements of events A and B respectively.



1
Expert's answer
2021-07-20T13:46:04-0400

i) If A and B are independent then "P(A \cap B) = P(A)P(B)". Then

"x(x + 0.2) = 0.15 \Rightarrow {x^2} + 0.2x - 0.15 = 0"

"D = 0.04 + 0.6 = 0.64"

"{x_1} = \frac{{ - 0.2 - 0.8}}{2} = - 0.5"

"{x_2} = \frac{{ - 0.2 + 0.8}}{2} = 0.3"

Since "P(A)>0" then "x=0.3".

Answer: "x=0.3"

ii)

"P(A) = x = 0.3"

"P(B) = x + 0.2 = 0.5"

Lets find

"P(A \cup B) = P(A) + P(B) - P(A \cap B)=0.3+0.5-0.15=0.65"

Lets find

"P({A^C}) = 1 - P(A) = 0.7"

"P({B^C}) = 1 - P(B) = 0.5"

Then

"P({A^C} \cap {B^C}) = P({A^C})P({B^C}) = 0.7 \cdot 0.5 = 0.35"

So

"P({A^C}\backslash {B^C}) = P({A^C}) - P({A^C} \cap {B^C}) = 0.7 - 0.35 = 0.35"

Answer: "P(A \cup B) =0.65" , "P({A^C}\backslash {B^C}) = 0.35"